### v4

# Ada-NAV: Adaptive Trajectory Length-Based Sample Efficient Policy Learning for Robotic Navigation

Abstract
Trajectory length stands as a crucial hyperparameter within reinforcement learning (RL) algorithms, significantly contributing to the sample inefficiency in robotics applications. Motivated by the pivotal role trajectory length plays in the training process, we introduce Ada-NAV, a novel adaptive trajectory length scheme designed to enhance the training sample efficiency of RL algorithms in robotic navigation tasks. Unlike traditional approaches that treat trajectory length as a fixed hyperparameter, we propose to dynamically adjust it based on the entropy of the underlying navigation policy. Interestingly, Ada-NAV can be applied to both existing on-policy and off-policy RL methods, which we demonstrate by empirically validating its efficacy on three popular RL methods: REINFORCE, Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC). We demonstrate through simulated and real-world robotic experiments that Ada-NAV outperforms conventional methods that employ constant or randomly sampled trajectory lengths. Specifically, for a fixed sample budget, Ada-NAV achieves an 18% increase in navigation success rate, a 20-38% reduction in navigation path length, and a 9.32% decrease in elevation costs. Furthermore, we showcase the versatility of Ada-NAV by integrating it with the Clearpath Husky robot, illustrating its applicability in complex outdoor environments.
I Introduction
Autonomous robotic navigation is ubiquitous in tasks such as search-and-rescue [1], disaster relief[2], and planetary exploration applications [3]. While traditional navigation methods have been popular among the community [4, 5, 6], reinforcement learning (RL) [7] has been recently successful in robotic navigation tasks [8]. RL-based methods offer advantages in terms of being mapless, having lower dependence on the sensor accuracy, and possessing the ability to work with raw sensor data [8, 9, 10]. However, the benefits of RL come with their own set of intricate challenges, most prominently the issue of exploration in the environment during training. Traditional RL methods often hinge on the availability of a dense reward structure [11], which may rarely be available in complex, unpredictable environments [12, 13, 14, 15]. These challenges make the existing RL methods sample inefficient during training, especially for robotics applications where rewards are sparse [16, 17], since the agent does not receive constant feedback on its progression towards the goal.
Figure 1: Robot navigation using policies trained with the REINFORCE policy gradient algorithm using four trajectory length schemes. We compare three fixed trajectory lengths, 200 (blue), 300 (green), and 300 with max entropy regularization (red), against Ada-NAV (ours) (yellow) in real-world environments. More details can be found in Sec. IV. The drawn trajectories are sample representations of the robot’s odometry data collected during each navigation task. Ada-NAV enhances sample efficiency of training (shown in the figure by the lower value of the orange bar plot) by adaptively varying trajectory length throughout training. Ada-NAV results in better navigation policies compared to baseline approaches and generates shorter and more successful test-time trajectories.
To improve the sample efficiency of RL for robotics navigation, common existing methods include reward-shaping [18] and learning from demonstrations [19, 20]. Reward-shaping methods use pseudo-rewards to encourage exploration in the environment [11, 21]. However, these methods require an expert to design intrinsic rewards, which can lead to performance-hindering bias. Furthermore, collecting expert-guided demonstrations for navigation tasks can be challenging and expensive to obtain [22].
In this work, we present Ada-NAV, a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods in the literature. Our key idea is to focus on adaptively changing a salient feature in most RL methods: trajectory length. We note that in policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training. For navigation, an interaction is any action the mobile robot takes and the resulting state in the environment and the reward it receives. It has been verified in the literature that short trajectory lengths can lead to inefficient learning, while long ones can be unreliable when in an unfamiliar environment [23]. Ada-NAV resolves this issue by having an adaptive trajectory length which is reliable and computationally tractable.
In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate Ada-NAV. To develop Ada-NAV, our novel approach is to start by connecting the trajectory length to the mixing time of the policy that generates the trajectory. The mixing time of a policy is the amount of time it takes for a robot to reach its limiting behavior under that policy. When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [24]. When mixing time is short, on the other hand, ensuring trajectory lengths are commensurate frees up effort for performing more rapid policy updates and environmental exploration. Unfortunately, mixing time is typically difficult to estimate in complex environments [25, 26], so proxy methods for determining trajectory length are needed.
Main Contributions: Ada-NAV increases sample efficiency by controlling the amount of data required to update the policy parameter, thus decreasing the number of training samples needed to learn an optimal policy for robot navigation tasks. Our specific contributions include:
  • We empirically establish a positive correlation between the policy entropy and the spectral gap of its induced Markov chain. We show that this correlation is present with various transition dynamics. This insight motivates our use of the changing policy entropy as a dynamic variable to control trajectory length.
  • We take advantage of this empirical insight by utilizing a monotonic mapping between policy entropy and trajectory length to facilitate exploration in environments with sparse rewards and serve as the foundation for our adaptive trajectory length scheme, Ada-NAV.
  • We present extensive empirical results using multiple RL methods, including REINFORCE [27] and Proximal Policy Optimization (PPO) [28] and Soft Actor-Critic (SAC) [29], to show that Ada-NAV improves the sample efficiency with both on-policy and off-policy methods. Our experiments in goal-reaching navigation tasks with both even and uneven terrains demonstrate that Ada-NAV outperforms constant trajectory length schemes, even those utilizing maximum entropy regularization, in terms of sample efficiency.
  • We evaluate the navigation performance of Ada-NAV in both simulated and real-world outdoor environments using a Clearpath Husky robot. We observe that, for a fixed sample budget, Ada-NAV leads to an 18% increase in navigation success rate, a 20-38% decrease in the navigation path length, and a 9.32% decrease in the elevation cost compared to the policies obtained by the other methods.
I-A Related Works
In this section we give a brief overview of the existing work related to RL for autonomous navigation and sparse rewards.
RL for Navigation: Motion planning and navigation has been well studied in robotics [4, 5, 6]. RL-based methods have been extensively used for robot navigations [30]. Methods such as DDPG [31], A3C [32], and DQN [33] have been used for outdoor navigation, while [34] learned a global map to find the shortest path out of a maze with RL. [35] used classical methods to create a hierarchical RL approach. For policy gradient methods, [36] demonstrated the superior performance of Soft Actor-Critic (SAC) over deep deterministic policy gradient (DDPG) in the navigation of mobile robots. In these policy gradient methods, it is well known that the gradient estimation requires lots of samples and the overall training procedure is sample inefficient.
Sparse Rewards: For model-free reinforcement learning, reward sparsity greatly affects policy search [37, 11] as a mobile robot will rarely receive any feedback to adjust its behavior. To produce more rewards which the algorithm can learn from, reward-shaping methods based on intrinsic rewards have been proposed. Methods in [38] and [21] modified the reward output to encourage exploration of the environment. [39] and [40] used imitation learning and demonstrations to overcome the challenges of sparse rewards. Curriculum learning has also been utilized to deal with sparse rewards in navigation. [41] divided the navigation task into a subgoal tree so the agent can learn from dense-to-sparse curriculum. [17] and [42] formulated a heavy-tailed policy to encourage exploration in sparse environments.
II Problem Formulation
II-A Reinforcement Learning and Policy Gradient Methods
We consider a Markov Decision Process (MDP) characterized by the tuple $\mathcal{M}:=(\mathcal{S},\mathcal{A},\mathbb{P},r)$, where $\mathcal{S}$ is the state space for the robot in the environment, $\mathcal{A}$ is the set of possible actions the robot can take, $\mathbb{P}$ is the transition probability kernel that determines the next state $s^{\prime}\in\mathcal{S}$ given $s\in\mathcal{S},a\in\mathcal{A}$ via $s^{\prime}\sim\mathbb{P}(\cdot|s,a)$, and $r:\mathcal{S}\times\mathcal{A}\to[0,r_{\max}]$ is the reward function. The mobile robot navigates through this environment using a (potentially stochastic) policy $\pi$ that provides a conditional probability distribution over actions $a$ given a state $s$. Given a policy $\pi$ and transition kernel $\mathbb{P}$, the transition dynamics of the Markov chain induced by $\pi$ over $\mathcal{M}$ is given by $\mathbb{P}_{\pi}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\mathbb{P}(s^{\prime}|s,a)\pi(a|s)$. The goal in the RL setting is to find a policy maximizing long-run reward over $\mathcal{M}$. To achieve this, one infinite-horizon objective typically considered in the RL literature is expected discounted reward given by
$J^{\gamma}(\pi)=\mathbb{E}_{\pi,\mu}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],$ (1)
where $\gamma\in(0,1)$ is some pre-specified discount factor and $\mu$ is a given initial start state distribution over $\mathcal{S}$. In (1), the expectation is over the policy and initial state distributions. In the discounted-reward settings, the state-action value function corresponding to a given policy $\pi$ is defined by $Q^{\gamma}_{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\right]$.
Policy Gradient Algorithm. To handle large, potentially continuous state and action spaces, we focus on the case where policy $\pi$ is parameterized by a vector $\theta\in\mathbb{R}^{d}$, where $d$ denotes the parameter dimension, leading to the notion of a parameterized policy $\pi_{\theta}$. We define $J^{\gamma}(\theta)=J^{\gamma}(\pi_{\theta})$. Then, following policy gradient theorem [43], we note that
$\nabla J^{\gamma}(\theta)=\mathbb{E}_{\pi_{\theta},\mu}\left[Q^{\gamma}_{\pi}(s,a)\nabla\log\pi_{\theta}(a|s)\right],$ (2)
which provides the gradient expression to perform gradient ascent on the objectives $J^{\gamma}(\theta)$. For the discounted setting, at time $t$ the policy gradient estimate is obtained by generating a $K\in\mathbb{N}$ length trajectory $\tau_{t}:=\{s_{t},a_{t},\ldots,s_{t+K-1}a_{t+K-1}\}$ using $\pi_{\theta_{t}}$, and then computing
$\widehat{\nabla J^{\gamma}(\theta_{t})}=\sum_{k=0}^{K-1}\gamma^{k}r(s_{t+k},a_{t+k})\nabla\log\pi_{\theta_{t}}(a_{t}|s_{t}).$ (3)
This expression is then used to perform a gradient update of the form $\theta_{t+1}=\theta_{t}+\alpha\widehat{\nabla J^{\gamma}(\theta_{t})}$, for some stepsize $\alpha>0$. In REINFORCE, increasing $K$ naturally improves the quality of the estimator given in Equation (3) [44], but at the cost of an increased number of samples. In reinforcement learning for robotic navigation, simulation data to train mobile agent can be expensive to obtain or generate. Thus, reducing the number of training sample needed without degrading navigation performance is crucial in mitigating operation costs. In the next section, we will explain the quantity mixing time and its importance to trajectory length selection.
II-B Mixing Time and Spectral Gap
In policy gradient algorithms an important issue is that, to obtain good estimates of the expectation (2), which is taken with respect to the stationary distribution induced over the underlying MDP by $\pi_{\theta}$, longer trajectories are often needed. The length of the trajectories is dictated by the mixing time – to be defined below – of the associated policy. Recent work on stochastic optimization [45] and policy gradient (PG) methods [24] has highlighted and clarified the connection between Markov chain mixing time and ideal trajectory length. To leverage this connection in our approach, we first present some relevant terminology and results.
We assume throughout that, for a given policy parameter $\theta$, the Markov chain induced by $\pi_{\theta}$ on the MDP $\mathcal{M}$ is irreducible and aperiodic, which is a common assumption in theoretical analyses of RL methods. First note that the ergodicity assumption guarantees that the Markov chain has a unique stationary distribution $\mu_{\theta}$. For an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},r)$ and policy $\pi_{\theta}$, the $\epsilon$-mixing time of $\pi_{\theta}$ is $\tau_{mix}^{\theta}(\epsilon):=\inf\{t:\sup_{s\in\mathcal{S}}\|\mathbb{P}_{\pi_{\theta}}^{t}(\cdot|s)-\mu_{\theta}(\cdot)\|_{TV}\leq\epsilon\}$. By convention, the mixing time is defined as $\tau_{mix}^{\theta}:=\tau_{mix}^{\theta}(1/4)$.
Though mixing times are useful in determining appropriate trajectory lengths, they are typically difficult to estimate in practice. Therefore, an alternative to estimating mixing time is estimating a lower bound. One such quantity that has been established to lower bound mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{*}_{\theta}$. As demonstrated in [46], the spectral gap and mixing time of a Markov chain bear the following relationship.
Theorem II.1
Let $\lambda_{\theta}\neq 1$ be an eigenvalue of the transition matrix $\mathbb{P}_{\pi_{\theta}}$. Then
$\tau_{mix}^{\theta}(\epsilon)\geq\left(\frac{1}{1-|\lambda_{\theta}|}-1\right)\log\left(\frac{1}{2\epsilon}\right).$ (4)
Taking $\epsilon=1/4$ and noticing that $\lambda^{*}_{\theta}$ gives the tightest lower bound, we have $\tau_{mix}^{\theta}\geq\left(1/(1-|\lambda^{*}_{\theta}|)-1\right)\log 2$. When an estimate of $\mathbb{P}_{\pi_{\theta}}$ is available, this provides a method for bounding its mixing time $\tau_{mix}^{\theta}$. Unfortunately, estimating $\mathbb{P}_{\pi_{\theta}}$ is not always practical in large and complex state and action spaces [25, 26]. Thus, a proxy for the spectral gap is crucial for navigation in complex environments. In the following section, we leverage this theorem to empirically establish a useful connection between policy entropy and spectral gap to adaptively change trajectory length with no need to estimate $\mathbb{P}_{\pi_{\theta}}$.
III Proposed Approach: Adaptive Trajectory-Based Policy Learning
Before presenting our main algorithm, Ada-NAV, we start by highlighting an interesting connection between a policy entropy and the spectral gap of its induced transition matrix.
III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap from the mixing time lower bound in Theorem 4. However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$. One possible workaround is to estimate the transition kernel, but this quickly becomes intractable due to the $\Omega(|\mathcal{S}||\mathcal{A}|)$ number of samples required for each estimate. Interestingly, one key insight we observe is that if the policy becomes more deterministic and thus the entropy decreases, the spectral gap of the induced transition matrix also decreases or stays the same, suggesting a monotonic relationship between mixing time and policy determinism.
Figure 2 illustrates this correlation with a 2D gridworld environment. We have a $25$-by-$25$ gridworld. Each plot in Figure 2 corresponds to a different fixed transition probability kernel $\mathbb{P}$. For each fixed kernel, we construct 100 tabular policies, each with a discrete action space of 5 elements to represent up, down, left, right, and stay. The tabular policies have various average entropies, ranging from maximum entropy with uniform probability over the action space for each state (end of x-axis on Figure 2) to completely deterministic for each state (beginning of x-axis) with a probability of $1$ of moving upwards. We define policy entropy in for tabular policies as the average Shannon entropy across all states, see Equation 5. For each policy, we multiply the policy with the fixed transition kernel to calculate the underlying Markov Chain and then estimate its spectral gap.
In this environment, if the agent chooses an action that would make it go off the edge, the agent will stay in the current state. Both transition kernels are deterministic: each state-action pair is mapped to a single next state a potential agent can occupy. One kernel represents an empty gridworld with no obstacles. The relationship between spectral gap and policy entropy is plotted in orange for this kernel. The other kernel is represents a gridworld with $4$ walls dividing it into $4$ equal regions. Please see Appendix for visualization. The plot corresponding to this kernel is blue in Figure 2.  [^1] Code to notebook to generate plot: https://github.com/Bridge00/spectral_gap_policy_entropy.git
To be clear, in these experiments, we have not trained any agent to navigate in the gridworld. We have a static environment where we manually designed transition kernel matrices and multiplied each one with 100 different tabular policies to estimate the underlying Markov Chains and their corresponding spectral gaps, a task normally intractable for more complex, realistic environments. For results displaying Ada-NAV’s effectiveness in increasing sample efficiency for a goal-reaching task in this environment and other gridworld environments, please see the Appendix.
Figure 2: Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over a $25$-by-$25$ gridworld environment with different fixed transition kernels. Policy entropy is a viable proxy for mixing time.
As previously mentioned, we define the policy entropy in the discrete action space case as average the Shannon entropy [47] over the state space:
$H(\pi_{\theta})=-\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mathbb{\pi}_{\theta}(a|s)\log\mathbb{\pi}_{\theta}(a|s).$ (5)
In the continuous action space case for the robotic navigation experiments in Section IV, we use the closed-form differential entropy for a multivariate Gaussian distribution,
$H(\pi_{\theta})=-\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}\frac{1}{2}\log((2\pi e)^{N}\det(\Sigma_{\theta}(s))).$ (6)
where $N$ is the action space dimension and $\Sigma_{\theta}(s)$ is the covariance matrix of the policy $\pi_{\theta}(\cdot|s)$. We note that there are no analytical equations to calculate mixing time, and calculating spectral gap requires knowledge of the transition dynamics $P(s^{\prime}|s,a)$ which restricts it to only model-based RL. Thus, using analytical Equations (5) and (6) as a proxy for mixing time is a more simple method that can be used for either model-free or model-based RL.
III-B Ada-NAV: Mapping Policy Entropy to Trajectory Length
We now present our novel adaptive trajectory length scheme for robotic navigation. The key to our approach is to select trajectory length based on the current policy entropy value. We perform this computation using a monotonic mapping from the range of possible policy entropy values to a range of possible trajectory length values. Let $H_{i}$ denote the maximum possible policy entropy value and $H_{c}$ the policy entropy of the current policy. In practice, we choose $H_{i}=H(\pi_{\theta_{0}})$ to be the entropy of the initial policy, since policy entropy tends to decrease during training. Let $t_{i}\in\mathbb{N}$ denote the user-specified initial trajectory lengths and let $\eta>0$ denote a user-specified sensitivity parameter that controls the rate of change of length (see the Appendix for specific values of $t_{i},\eta$ used). For our experiments, we determine the trajectory length corresponding to the current policy entropy using the mapping $f$ below:
$t_{c}=f^{H_{i}}_{t_{i},\eta}(H_{c})=\text{round}\left(1-\frac{H_{c}}{H_{i}}(1-\frac{t_{i}}{\eta})\right),$ (7)
where the rounding operation is to ensure that $t_{c}$ is always an integer. In the Appendix, we utilize other monotonic transformation to map entropy to trajectory length in the form of an exponential function. Furthermore, in our experiments, we initialize our policies with a Gaussian distribution. To ensure, that $H_{c}$ never increases too much so that $t_{c}$ becomes 00, one can initialize with a uniform distribution so that $H_{i}$ is the maximum possible entropy. Algorithm 1 summarizes our Ada-NAV methodology for adaptive trajectory length PG methods. Note that any on-policy PG method can used to carry out the policy gradient update in line 3. The training is divided into episodes where the start location and goal are constant between episodes. During each episode, a gradient update occurs after every trajectory. The amount of samples in each trajectory is the trajectory length. The episode ends when goal is reached or when a set number of interactions have occurred. For clarification, given an episode, the trajectory length is fixed. At the start of the new episode is when the trajectory length is adjusted based on policy entropy.
1:  Initialize: policy parameter $\theta_{0}$, policy entropy $H_{i}$, trajectory length $t_{i}$, current trajectory length $t_{c}=t_{i}$, maximum trajectory length $t_{d}$, number of episodes $E$, maximum number $K$ of iterations per episode, sensitivity parameter $\eta$, $t=0$
2:  for $e=0,\ldots,E-1$ do
3:     calculate $H_{c}$ using 5 for discrete actions, 6 for continuous actions
4:     calculate $t_{c}$ using 7
5:     $k=0$
6:     while $k<K$ do
[…]
8:        Perform gradient update using $\mathcal{T}$ to obtain $\theta_{t+1}$
9:        $k=k+t_{c}$
10:        $t=t+1$
11:     end while
12:  end for
13:  Return: $\pi_{\theta_{t}}$
Algorithm 1 Policy Learning with Ada-NAV
IV Experiments and Results
This section provides a detailed empirical study and comparison of the proposed adaptive trajectory methods in Algorithm 1.
Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator. For REINFORCE, PPO, and SAC, using Ada-NAV as the trajectory length scheme leads to a navigation policy with a higher cumulative reward return than using constant trajectory lengths. The drawn trajectories are sample representations of the robot’s odometry data collected during each navigation task. The trajectories pictured are representative of all three algorithms, since we found that each trajectory-length scheme leads to similar navigation behavior regardless of which RL algorithm is used. See Figures 4 and 6 for training details.
Figure 4: [TOP] Learning curves during training and [BOTTOM] sample complexity savings during navigation tasks for the even terrain shown on the left of Figure 3. Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, Ada-NAV converges to a higher cumulative reward while expending less of the total budget of $\sim 4.5e4$ samples. We ran $8$ to $10$ independent replications for each of the 12 experiments. Although fixed trajectory length 200 performs comparably to Ada-NAV in terms of sample complexity, it achieves lower reward in all cases.
Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained using Ada-NAV vs. fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$. We observe that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training. In contrast, Ada-NAV policies successfully complete the navigation tasks after training with the same sample budget. Further, the experiments with a real Clearpath Husky robot validate that our method can be transferred onto real robotic systems without significant performance degradation. The drawn trajectories are sample representations of the robot’s odometry data collected during each task.
RL Algorithms. We choose three RL algorithms to integrate Ada-NAV into for improved sample efficiency: REINFORCE, PPO, and SAC. As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that Ada-NAV can improve the sample efficiency of various classes of policy gradient algorithms.
Baselines. Since our proposed approach relies on adaptive trajectory length, for each algorithm we test Ada-NAV with, we compare it against the same algorithm with fixed trajectory lengths. We also compare with fixed trajectory lengths with max-entropy regularizer as that also encourages exploration.
We also incorporate a classical navigation algorithm Ego-graph [48], to highlight the navigation performance of our method under challenging terrain conditions. Ego-graph is a baseline method in [48] that utilizes robot-centric elevation data to obtain the actions that minimize the elevation gradient cost.
Figure 6: [TOP] Learning curves during training and [BOTTOM] sample complexity savings during navigation tasks for the uneven terrain shown on the right of Figure 3. Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, Ada-NAV converges to a higher cumulative reward while expending less of the total budget of $\sim 4.5e4$ samples. We ran $8$ to $10$ independent replications for each of the 12 experiments. Although fixed trajectory length 200 performs comparably to Ada-NAV in terms of sample complexity, it achieves lower reward in all cases.
Environments: We utilize two environments in a unity-based outdoor robot simulator. The policies for the outdoor robot simulator are based on the REINFORCE algorithm [27]. Finally, we deploy navigation policies trained for a fixed sample budget using the robot simulator on a real Clearpath Husky robot. We present the details of the training environments below:
Even and Uneven Terrain Navigation in Robotic Simulations: We use a Unity-based outdoor simulator with a Clearpath Husky robot model to train policies for two navigation tasks: 1. Goal Reaching, 2. Uneven terrain navigation using the Reinforce algorithms. The Unity simulator includes diverse terrains and elevations for training and testing. We restrict ourselves to a $100m\times 100m$ region during training. We used identical policy distribution parameters and neural network architectures with all four algorithms during training and evaluation. The policies are implemented using Pytorch in to train in the simulator equipped with a Clearpath Husky robot model, and a ROS Melodic platform. The training and simulations are executed on a workstation with an Intel Xeon 3.6 GHz CPU and an Nvidia Titan GPU.
Our agent is a differential drive robot with a two-dimensional continuous action space $a=(v,\omega)$ (i.e. linear and angular velocities). The state space varies for the two navigation tasks and is obtained in real time from the simulated robot’s odometry and IMU sensors. The state inputs used in the two scenarios are, The simulator includes even and uneven outdoor terrains and a differential drive robot model with a two-dimensional continuous action space $a=(v,\omega)$ (i.e. linear and angular velocities). We perform two navigation tasks: 1. Even terrain, 2. Uneven terrain navigation using the robot’s local sensor data as state observations. The state space varies for the two navigation tasks and is obtained in real-time from the simulated robot’s odometry and IMU sensors. The state inputs used in the two scenarios are,
  • $d_{goal}\in[0,1]$ Normalized distance (w.r.t the initial straight-line distance) between the robot and its goal;
  • $\alpha_{goal}\in[0,\pi]$ Angle between the robot’s heading direction and the goal;
  • $(\phi,\theta)\in[0,\pi]$ Roll and pitch angle of the robot respectively.
Inspired by [17], we define three reward functions to formulate the sparse rewards to train our navigation policies: Goal distance reward $r_{dist}$, Goal heading reward $r_{head}$, and Elevation reward $r_{elev}$ to formulate the sparse rewards to train our navigation policies. Hence,
$r_{dist}=\frac{\beta}{2}\mathcal{N}(\frac{d_{goal}}{2},\sigma^{2})+\beta\mathcal{N}(0,\sigma^{2}),$ (8)
$r_{head}=\mathbb{I}_{\{|\alpha_{goal}|\leq\pi/3\}},$ (9)
where, $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter. We set the variance $\sigma=0.2$ to ensure that the $r_{dist}$ only includes two narrow reward peaks near the goal and half-way from the goal. Similarly, $r_{elev}$ is defined as $r_{elev}=\mathbb{I}_{\{|\phi|\geq\pi/6\}}\cup\mathbb{I}_{\{|\theta|\geq\pi/6\}}$. We observe that our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature[31].
(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position are given at each episode. The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{goal}+r_{head}$.
(b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment. The state space is $s=[d_{goal},\alpha_{goal},\phi,\theta]$. The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs. The overall reward formulation $r_{uneven}=r_{goal}+r_{head}+r_{elev}$.
(c) Real-world Experiments: The real and simulated testing were conducted in different outdoor scenarios. Particularly, the real-world environments are reasonably different in terms of terrain structure (i.e. elevation gradient) and surface properties (i.e grass and tiny gravel regions with different levels of friction) which are not included or modeled in the simulator (e.g. elevation gain in the simulator is up to $\sim 4m$, however, we restricted it to $\sim 1.5m$ elevation gain in the real world for robot’s safety). Our Clearpath Husky robot is equipped with a VLP16 LiDAR and a laptop with an Intel i9 CPU and an Nvidia RTX 2080 GPU.
[…]
tabular
Terrain Algorithm SR PL EC
(%$\uparrow$ (# m)$\downarrow$ (m)$\downarrow$
Even Reinforce: traj. length = 300 + max entropy 58 22.47 -
Reinforce: traj. length = 200 32 25.58 -
Reinforce: traj. length = 300 65 18.75 -
Ada-NAV(Ours) 84 11.62 -
Uneven Ego-graph [48] 54 20.36 2.189
Reinforce: traj. length = 300 + max entropy 28 41.32 3.163
Reinforce: traj. length = 200 8 9.25 3.116
Reinforce: traj. length = 300 31 32.44 2.236
Ada-NAV(Ours) 72 16.28 1.985
Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks at a fixed sample budget. SR denotes success rate, PL denotes average path length, and EC denotes elevation cost. Each method is tested for 100 trials in the outdoor simulator and the average values are reported in the table. Test environments are different from training environments.
IV-A Evaluation Metrics
We use the following metrics to evaluate the quality of the navigation paths generated with a fixed sample budget:
  • Success rate: The percentage of successful goal-reaching attempts without robot flip-overs (especially in uneven terrain) out of the total number of trials.
  • Average path length: The average length of the path taken to reach a goal located 10 meters away from the robot’s starting position.
  • Elevation cost: The total elevation gradient experienced by the robot during a trial (i.e., if $z_{r}$ is the vector that includes the gradient of the vertical motions of the robot along a path, the elevation cost is given by $||\nabla z_{r}||$).
IV-B Results and Discussions
Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, Ada-NAV achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths. Significantly, despite providing comparable returns to Ada-NAV, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig. 4 and 6). Please see the Appendix for similar results in 2D gridworld environments with actor-critic algorithm[49].
Navigation Performance: We compare our method’s navigation performance qualitatively in Fig. 5 and quantitatively in Table I. To highlight the importance of sample efficiency during training, we use policies trained for a fixed sample budget to perform navigation tasks in even and uneven simulated terrains. We observe that, except Ada-NAV, the policies cannot consistently complete the navigation tasks, resulting in significantly lower success rates. Even successful trials of these policies lead to significantly longer paths with several loops instead of heading toward the goal. In contrast, Ada-NAV results in relatively straight and consistent paths toward the goals. We further observe that uneven terrain navigation is a comparatively challenging task even for traditional methods such as Ego-graph[48]. Hence, the success rates are relatively low due to the robot flip-overs that occurred during trials. However, we observe that Ada-NAV leads to higher success rates even compared to baseline methods such as ego-graph. We further observe that Ada-NAV results in comparatively lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions. Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot. We observe that our Ada-NAV even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies demonstrate inconsistent navigation performance similar to the simulations.
V Conclusions, Limitations, and Future Works
In this work, we propose an adaptive trajectory scheme, called Ada-NAV, which encourages exploration and induces training sample efficiency for goal-reaching navigation in the sparse reward setting with RL-based methods. We do so by empirically establishing a positive correlation between a policy’s entropy and the spectral gap of its induced Markov Chain. We show, through experiments in navigation tasks, that Ada-NAV induces more sample-efficient training than both fixed and random trajectory length schemes without the need for tuning. In experiments on real robots, Ada-NAV-trained policies generally achieve better success rates and shorter path lengths than policies trained using constant and random trajectory lengths for the same sample budget.
An important limitation of the present work is the restricted scope of our experimental investigation of the connection between policy entropy and the induced spectral gap. We considered only simple gridworld environments for the purposes of this paper, yet this important connection merits further investigation and validation on more sophisticated challenges, such as the MuJoCo suite of RL problems [50]. During real-world experiments, we observe that all the navigation policies often lead to high motor vibrations due to dynamically infeasible velocities provided as the action. Hence, further investigation is required to generate dynamically feasible and smooth actions for a better sim-to-real transfer.

### v7

@@
-Ada-NAV: Adaptive Trajectory Length-Based Sample Efficient Policy Learning for Robotic Navigation
+CCE: Sample Efficient Sparse Reward Policy Learning for Robotic Navigation via Confidence-Controlled Exploration
 Abstract
-Trajectory length stands as a crucial hyperparameter within reinforcement learning (RL) algorithms, significantly contributing to the sample inefficiency in robotics applications.
-Motivated by the pivotal role trajectory length plays in the training process, we introduce Ada-NAV, a novel adaptive trajectory length scheme designed to enhance the training sample efficiency of RL algorithms in robotic navigation tasks.
-Unlike traditional approaches that treat trajectory length as a fixed hyperparameter, we propose to dynamically adjust it based on the entropy of the underlying navigation policy.
-Interestingly, Ada-NAV can be applied to both existing on-policy and off-policy RL methods, which we demonstrate by empirically validating its efficacy on three popular RL methods: REINFORCE, Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC).
-We demonstrate through simulated and real-world robotic experiments that Ada-NAV outperforms conventional methods that employ constant or randomly sampled trajectory lengths.
-Specifically, for a fixed sample budget, Ada-NAV achieves an 18% increase in navigation success rate, a 20-38% reduction in navigation path length, and a 9.32% decrease in elevation costs.
-Furthermore, we showcase the versatility of Ada-NAV by integrating it with the Clearpath Husky robot, illustrating its applicability in complex outdoor environments.
+We introduce Confidence-Controlled Exploration (CCE), a novel exploration scheme designed to enhance the training sample efficiency of reinforcement learning (RL) algorithms for sparse reward settings such as robot navigation.
+Sparse rewards are common in RL and convenient to design and implement, but typically hard to deal with due to the challenges of exploration.
+Existing methods deploy regularization-based methods to deal with the exploration challenges.
+However, it is hard to characterize the balance between exploration and exploitation because regularization modifies the reward function itself, hence changing the objective we are optimizing for.
+In contrast to regularization-based approaches in the existing literature, our approach, CCE, is based on a novel relationship we provide between gradient estimation and policy entropy.
+CCE dynamically adjusts the number of samples of the gradient update used during training to control exploration.
+Interestingly, CCE can be applied to both existing on-policy and off-policy RL methods, which we demonstrate by empirically validating its efficacy on three popular RL methods: REINFORCE, Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC) for goal-reaching robotic navigation tasks.
+We demonstrate through simulated and real-world experiments that CCE outperforms conventional methods that employ constant trajectory lengths and entropy regularization when constraining the sample budget.
+For a fixed sample budget, CCE achieves an 18% increase in navigation success rate, a 20-38% reduction in navigation path length, and a 9.32% decrease in elevation costs.
+Furthermore, we showcase the versatility of CCE by integrating it with the Clearpath Husky robot, illustrating its applicability in complex outdoor environments.
 I Introduction
-Autonomous robotic navigation is ubiquitous in tasks such as search-and-rescue [1], disaster relief[2], and planetary exploration applications [3].
-While traditional navigation methods have been popular among the community [4, 5, 6], reinforcement learning (RL) [7] has been recently successful in robotic navigation tasks [8].
-RL-based methods offer advantages in terms of being mapless, having lower dependence on the sensor accuracy, and possessing the ability to work with raw sensor data [8, 9, 10].
-However, the benefits of RL come with their own set of intricate challenges, most prominently the issue of exploration in the environment during training.
-Traditional RL methods often hinge on the availability of a dense reward structure [11], which may rarely be available in complex, unpredictable environments [12, 13, 14, 15].
-These challenges make the existing RL methods sample inefficient during training, especially for robotics applications where rewards are sparse [16, 17], since the agent does not receive constant feedback on its progression towards the goal.
+Reinforcement learning (RL) [1] is an important component of learning in robotic systems in a wide variety of complex domains, including search-and-rescue [2], disaster relief [3], and planetary exploration [4] applications.
+However, traditional RL methods often hinge on the availability of a dense reward structure [5], which may rarely be available in complex, unpredictable environments [6, 7, 8, 9].
+Thus, reward function sparsity [10] frequently arise when applying RL methods to robotics problems, such as robotic navigation, due to their simple to implement and does not require expert knowledge to design like dense reward structures.
+However, this simplicity comes with a sparse feedback signal to learn from, thus leading to a limited exploration of the environment, and ultimately, to a suboptimal policy [11].
+Existing literature in RL for robotics in sparse rewards has utilized methods such as entropy regularization [12], reward shaping [13, 14, 15, 16], curiosity-driven exploration [5, 17].
+However, these methods are regularization-based that change the reward optimization objective which can lead to instability in learning [6].
+Another major challenge of existing works comes from the sample inefficiency of exploration in sparse rewards.
+This issue is significant in robotic systems, due to the simulation-intensive nature of RL methods and the computational expense of realistic robotics simulation platforms [18].
+Thus, RL training typically requires finding an optimal policy in a constrained sample budget.
+In this work, we diverge from previous exploration paradigms to respect the practicality of limited training samples.
+We do so by controlling exploration, rather than encouraging it based on a novel connection between the uncertainty, or confidence, of a policy and policy gradient estimation.
 Figure 1: Robot navigation using policies trained with the REINFORCE policy gradient algorithm using four trajectory length schemes.
-We compare three fixed trajectory lengths, 200 (blue), 300 (green), and 300 with max entropy regularization (red), against Ada-NAV (ours) (yellow) in real-world environments.
+We compare three fixed trajectory lengths, 200 (blue), 300 (green), and 300 with max entropy regularization (red), against CCE (ours) (yellow) in real-world environments.
 More details can be found in Sec.
 IV.
-The drawn trajectories are sample representations of the robot’s odometry data collected during each navigation task.
-Ada-NAV enhances sample efficiency of training (shown in the figure by the lower value of the orange bar plot) by adaptively varying trajectory length throughout training.
-Ada-NAV results in better navigation policies compared to baseline approaches and generates shorter and more successful test-time trajectories.
-To improve the sample efficiency of RL for robotics navigation, common existing methods include reward-shaping [18] and learning from demonstrations [19, 20].
-Reward-shaping methods use pseudo-rewards to encourage exploration in the environment [11, 21].
-However, these methods require an expert to design intrinsic rewards, which can lead to performance-hindering bias.
-Furthermore, collecting expert-guided demonstrations for navigation tasks can be challenging and expensive to obtain [22].
-In this work, we present Ada-NAV, a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods in the literature.
-Our key idea is to focus on adaptively changing a salient feature in most RL methods: trajectory length.
-We note that in policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training.
+The drawn trajectories are sample representations of the robot’s odometry data collected.
+CCE enhances training sample efficiency (shown in the figure by the lower value of the orange bar plot) by adaptively varying trajectory length.
+CCE results in better navigation policies compared to baseline approaches and generates shorter and more successful test-time trajectories.
+Main Contributions: We present Confidence-Controlled Exploration (CCE), a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods.
+We introduce using the policy confidence, measured by the average conditional policy entropy over all states to control exploration.
+We focus on balancing exploration with the policy entropy by adaptively changing a salient feature in most RL methods: trajectory length.
+In policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training.
+Our novel perspective is that exploration can be adaptively controlled by the number of samples needed to estimate the gradient.
 For navigation, an interaction is any action the mobile robot takes and the resulting state in the environment and the reward it receives.
-It has been verified in the literature that short trajectory lengths can lead to inefficient learning, while long ones can be unreliable when in an unfamiliar environment [23].
-Ada-NAV resolves this issue by having an adaptive trajectory length which is reliable and computationally tractable.
-In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate Ada-NAV.
-To develop Ada-NAV, our novel approach is to start by connecting the trajectory length to the mixing time of the policy that generates the trajectory.
+To design CCE, our novel approach connects the need for the policy to explore to the mixing time of the policy.
 The mixing time of a policy is the amount of time it takes for a robot to reach its limiting behavior under that policy.
 When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [19].
-When mixing time is short, on the other hand, ensuring trajectory lengths are commensurate frees up effort for performing more rapid policy updates and environmental exploration.
+When mixing time is short, on the other hand, ensuring trajectory lengths are proportional frees up effort for performing more rapid policy updates and environmental exploration.
 Unfortunately, mixing time is typically difficult to estimate in complex environments [20, 21], so proxy methods for determining trajectory length are needed.
-Main Contributions: Ada-NAV increases sample efficiency by controlling the amount of data required to update the policy parameter, thus decreasing the number of training samples needed to learn an optimal policy for robot navigation tasks.
-Our specific contributions include:
+In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate CCE.
+We increase the sample efficiency by adaptively tuning the trajectory length required to update the policy parameter, thus decreasing the number of samples needed to learn an optimal policy for navigation tasks.
+Our results include:
   • We empirically establish a positive correlation between the policy entropy and the spectral gap of its induced Markov chain.
 We show that this correlation is present with various transition dynamics.
-This insight motivates our use of the changing policy entropy as a dynamic variable to control trajectory length.
-  • We take advantage of this empirical insight by utilizing a monotonic mapping between policy entropy and trajectory length to facilitate exploration in environments with sparse rewards and serve as the foundation for our adaptive trajectory length scheme, Ada-NAV.
-  • We present extensive empirical results using multiple RL methods, including REINFORCE [27] and Proximal Policy Optimization (PPO) [28] and Soft Actor-Critic (SAC) [29], to show that Ada-NAV improves the sample efficiency with both on-policy and off-policy methods.
-Our experiments in goal-reaching navigation tasks with both even and uneven terrains demonstrate that Ada-NAV outperforms constant trajectory length schemes, even those utilizing maximum entropy regularization, in terms of sample efficiency.
-  • We evaluate the navigation performance of Ada-NAV in both simulated and real-world outdoor environments using a Clearpath Husky robot.
-We observe that, for a fixed sample budget, Ada-NAV leads to an 18% increase in navigation success rate, a 20-38% decrease in the navigation path length, and a 9.32% decrease in the elevation cost compared to the policies obtained by the other methods.
+This insight motivates our use of the changing policy entropy to control the amount of exploration during training.
+  • We take advantage of this empirical insight by utilizing a monotonic mapping between policy entropy and trajectory length to facilitate exploration with sparse rewards and serve as the foundation for CCE.
+  • We present extensive empirical results using multiple RL methods, including REINFORCE [22] and Proximal Policy Optimization (PPO) [23] and Soft Actor-Critic (SAC) [12], to show that CCE improves the sample efficiency with both on-policy and off-policy methods.
+Furthermore, by showing results on SAC, we show that CCE also improves sample efficiency for algorithms that employ entropy regularization.
+Our experiments in goal-reaching navigation tasks with both even and uneven terrains demonstrate that CCE outperforms constant trajectory length and exploration encouragement methods, with constrained samples.
+  • We evaluate the navigation performance of CCE in both simulated and real-world outdoor environments using a Clearpath Husky robot.
+We observe that, for a fixed, limited sample budget, CCE leads to an 18% increase in navigation success rate, a 20-38% decrease in the navigation path length, and a 9.32% decrease in the elevation cost compared to the policies from baselines.
 I-A Related Works
-In this section we give a brief overview of the existing work related to RL for autonomous navigation and sparse rewards.
-RL for Navigation: Motion planning and navigation has been well studied in robotics [4, 5, 6].
-RL-based methods have been extensively used for robot navigations [30].
-Methods such as DDPG [31], A3C [32], and DQN [33] have been used for outdoor navigation, while [34] learned a global map to find the shortest path out of a maze with RL.
-[35] used classical methods to create a hierarchical RL approach.
-For policy gradient methods, [36] demonstrated the superior performance of Soft Actor-Critic (SAC) over deep deterministic policy gradient (DDPG) in the navigation of mobile robots.
-In these policy gradient methods, it is well known that the gradient estimation requires lots of samples and the overall training procedure is sample inefficient.
 Sparse Rewards: For model-free reinforcement learning, reward sparsity greatly affects policy search [24, 5] as a mobile robot will rarely receive any feedback to adjust its behavior.
@@
 Methods in [13] and [25] modified the reward output to encourage exploration of the environment.
-[39] and [40] used imitation learning and demonstrations to overcome the challenges of sparse rewards.
-Curriculum learning has also been utilized to deal with sparse rewards in navigation.
-[41] divided the navigation task into a subgoal tree so the agent can learn from dense-to-sparse curriculum.
-[17] and [42] formulated a heavy-tailed policy to encourage exploration in sparse environments.
+Furthermore, [26] and [27] formulated a heavy-tailed policy to encourage exploration in sparse environments.
+Prior work used imitation learning and demonstration [15, 28] and curriculum learning [29] to overcome the challenges of sparse rewards.
+Exploration and Exploitation in RL and Navigation: To balance exploration and exploitation, prior work has used the uncertainty of a surrogate cost function estimation from the visual field of view [30] and a world model in model-based RL [31, 32].
+In motion planning, [33] introduced a planner that balances exploration and exploitation using explicit information on the environmental information.
+By connecting exploration to policy gradient estimation, our approach, CEE, is agnostic to data modality and does not require a world model or any explicit information on the given environment.
+Exploration-exploitation for multi-robot systems was considered in [34, 35, 36].
+RL for Navigation: Motion planning and navigation has been well studied in robotics [37, 38, 39].
+RL-based methods have been extensively used for robot navigations [40].
+Methods such as in [41] used DRL to find reliable paths for outdoor navigation, while [42] learned a global map to find the shortest path out of a maze with RL.
+In [43], classical methods were used to create a hierarchical RL approach.
 II Problem Formulation
@@
 The mobile robot navigates through this environment using a (potentially stochastic) policy $\pi$ that provides a conditional probability distribution over actions $a$ given a state $s$.
-Given a policy $\pi$ and transition kernel $\mathbb{P}$, the transition dynamics of the Markov chain induced by $\pi$ over $\mathcal{M}$ is given by $\mathbb{P}_{\pi}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\mathbb{P}(s^{\prime}|s,a)\pi(a|s)$.
+Given a policy $\pi$ and transition kernel $\mathbb{P}$, the transition dynamics of the ergodic Markov chain induced by $\pi$ over $\mathcal{M}$ is given by $\mathbb{P}_{\pi}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\mathbb{P}(s^{\prime}|s,a)\pi(a|s)$.
 The goal in the RL setting is to find a policy maximizing long-run reward over $\mathcal{M}$.
-To achieve this, one infinite-horizon objective typically considered in the RL literature is expected discounted reward given by
-$J^{\gamma}(\pi)=\mathbb{E}_{\pi,\mu}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],$ (1)
-where $\gamma\in(0,1)$ is some pre-specified discount factor and $\mu$ is a given initial start state distribution over $\mathcal{S}$.
+To achieve this, let $\mu$ be the unique stationary distribution to the ergodic Markov Chain.
+Thus, one infinite-horizon objective typically considered in the RL literature is the expected discounted reward given by
+$J(\pi)=\mathbb{E}_{s\sim\mu,a\sim\pi}\left[\sum_{t=0}^{\infty}r(s_{t},a_{t})\right],$ (1)
 In (1), the expectation is over the policy and initial state distributions.
-In the discounted-reward settings, the state-action value function corresponding to a given policy $\pi$ is defined by $Q^{\gamma}_{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\right]$.
+In the average-reward settings, the state-action value function corresponding to a given policy $\pi$ is defined by $Q_{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}r(s_{t},a_{t})-J(\pi)\ |\ s_{0}=s,a_{0}=a\right]$.
 Policy Gradient Algorithm.
 To handle large, potentially continuous state and action spaces, we focus on the case where policy $\pi$ is parameterized by a vector $\theta\in\mathbb{R}^{d}$, where $d$ denotes the parameter dimension, leading to the notion of a parameterized policy $\pi_{\theta}$.
-We define $J^{\gamma}(\theta)=J^{\gamma}(\pi_{\theta})$.
+We define $J(\theta)=J(\pi_{\theta})$.
 Then, following policy gradient theorem [44], we note that
-$\nabla J^{\gamma}(\theta)=\mathbb{E}_{\pi_{\theta},\mu}\left[Q^{\gamma}_{\pi}(s,a)\nabla\log\pi_{\theta}(a|s)\right],$ (2)
-which provides the gradient expression to perform gradient ascent on the objectives $J^{\gamma}(\theta)$.
-For the discounted setting, at time $t$ the policy gradient estimate is obtained by generating a $K\in\mathbb{N}$ length trajectory $\tau_{t}:=\{s_{t},a_{t},\ldots,s_{t+K-1}a_{t+K-1}\}$ using $\pi_{\theta_{t}}$, and then computing
-$\widehat{\nabla J^{\gamma}(\theta_{t})}=\sum_{k=0}^{K-1}\gamma^{k}r(s_{t+k},a_{t+k})\nabla\log\pi_{\theta_{t}}(a_{t}|s_{t}).$ (3)
-This expression is then used to perform a gradient update of the form $\theta_{t+1}=\theta_{t}+\alpha\widehat{\nabla J^{\gamma}(\theta_{t})}$, for some stepsize $\alpha>0$.
-In REINFORCE, increasing $K$ naturally improves the quality of the estimator given in Equation (3) [44], but at the cost of an increased number of samples.
-In reinforcement learning for robotic navigation, simulation data to train mobile agent can be expensive to obtain or generate.
-Thus, reducing the number of training sample needed without degrading navigation performance is crucial in mitigating operation costs.
-In the next section, we will explain the quantity mixing time and its importance to trajectory length selection.
+$\nabla J(\theta)=\mathbb{E}_{\pi_{\theta},\mu}\left[Q_{\pi}(s,a)\nabla\log\pi_{\theta}(a|s)\right],$ (2)
+which provides the gradient expression to perform gradient ascent on the objectives $J(\theta)$.
+At update $u$ the policy gradient estimate is obtained by generating a $K\in\mathbb{N}$ length trajectory $\tau_{t}:=\{s_{0},a_{0},\ldots,s_{K-1}a_{K-1}\}$ using $\pi_{\theta_{u}}$, and then computing
+$\widehat{\nabla J(\theta_{u})}=\frac{1}{K-1}\sum_{k=0}^{K-1}r(s_{k},a_{k})\nabla\log\pi_{\theta_{t}}(a_{k}|s_{k}).$ (3)
+This expression is then used to perform a gradient update of the form $\theta_{u+1}=\theta_{u}+\eta\widehat{\nabla J^{(}\theta_{u})}$, for some stepsize $\eta>0$.
+In REINFORCE, increasing $K$, and therefore the overall number of samples naturally improves the quality of the estimator given in Equation (3) [45].
+However, in reinforcement learning for robotic navigation, simulation data to train mobile agent can be expensive to obtain or generate.
+Thus, reducing the number of training samples needed without degrading navigation performance is crucial in mitigating operation costs.
+In the next section, we will explain mixing time and its importance to gradient estimation.
 II-B Mixing Time and Spectral Gap
-In policy gradient algorithms an important issue is that, to obtain good estimates of the expectation (2), which is taken with respect to the stationary distribution induced over the underlying MDP by $\pi_{\theta}$, longer trajectories are often needed.
+In policy gradient algorithms an important issue is that, to obtain good estimates of the expectation in (2), which is taken with respect to the stationary distribution, $\mu_{\theta}$, induced over the underlying MDP by $\pi_{\theta}$, longer trajectories are often needed.
 The length of the trajectories is dictated by the mixing time – to be defined below – of the associated policy.
-Recent work on stochastic optimization [45] and policy gradient (PG) methods [24] has highlighted and clarified the connection between Markov chain mixing time and ideal trajectory length.
-To leverage this connection in our approach, we first present some relevant terminology and results.
-We assume throughout that, for a given policy parameter $\theta$, the Markov chain induced by $\pi_{\theta}$ on the MDP $\mathcal{M}$ is irreducible and aperiodic, which is a common assumption in theoretical analyses of RL methods.
+Recent work on stochastic optimization [46] and policy gradient (PG) methods [19] has highlighted and clarified the connection between Markov chain mixing time and ideal trajectory length, which we will later use to connect confidence to exploration.
+We assume throughout that, for a given policy parameter $\theta$, the Markov chain induced by $\pi_{\theta}$ on the MDP $\mathcal{M}$ is ergodic, which is a common assumption in theoretical analyses of RL methods [19, 47, 46, 48].
 First note that the ergodicity assumption guarantees that the Markov chain has a unique stationary distribution $\mu_{\theta}$.
 For an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathbb{P},r)$ and policy $\pi_{\theta}$, the $\epsilon$-mixing time of $\pi_{\theta}$ is $\tau_{mix}^{\theta}(\epsilon):=\inf\{t:\sup_{s\in\mathcal{S}}\|\mathbb{P}_{\pi_{\theta}}^{t}(\cdot|s)-\mu_{\theta}(\cdot)\|_{TV}\leq\epsilon\}$.
-By convention, the mixing time is defined as $\tau_{mix}^{\theta}:=\tau_{mix}^{\theta}(1/4)$.
-Though mixing times are useful in determining appropriate trajectory lengths, they are typically difficult to estimate in practice.
-Therefore, an alternative to estimating mixing time is estimating a lower bound.
-One such quantity that has been established to lower bound mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{*}_{\theta}$.
-As demonstrated in [46], the spectral gap and mixing time of a Markov chain bear the following relationship.
-Theorem II.1
-Let $\lambda_{\theta}\neq 1$ be an eigenvalue of the transition matrix $\mathbb{P}_{\pi_{\theta}}$. Then
-$\tau_{mix}^{\theta}(\epsilon)\geq\left(\frac{1}{1-|\lambda_{\theta}|}-1\right)\log\left(\frac{1}{2\epsilon}\right).$ (4)
-Taking $\epsilon=1/4$ and noticing that $\lambda^{*}_{\theta}$ gives the tightest lower bound, we have $\tau_{mix}^{\theta}\geq\left(1/(1-|\lambda^{*}_{\theta}|)-1\right)\log 2$.
-When an estimate of $\mathbb{P}_{\pi_{\theta}}$ is available, this provides a method for bounding its mixing time $\tau_{mix}^{\theta}$.
-Unfortunately, estimating $\mathbb{P}_{\pi_{\theta}}$ is not always practical in large and complex state and action spaces [25, 26].
+By convention, the mixing time is defined as $\tau_{mix}^{\theta}:=\tau_{mix}^{\theta}(1/4)$ [47].
+Though mixing times are useful in determining trajectory lengths, they are difficult to estimate in practice.
+Therefore, an alternative is estimating a lower bound.
+As described in [49], one such lower bound of mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{2}_{\theta}$:
+When an estimate of $\mathbb{P}_{\pi_{\theta}}$ is available, this provides a method for bounding its mixing time.
+Unfortunately, estimating $\mathbb{P}_{\pi_{\theta}}$ is impractical in large and complex state and action spaces [20, 21].
 Thus, a proxy for the spectral gap is crucial for navigation in complex environments.
-In the following section, we leverage this theorem to empirically establish a useful connection between policy entropy and spectral gap to adaptively change trajectory length with no need to estimate $\mathbb{P}_{\pi_{\theta}}$.
-III Proposed Approach: Adaptive Trajectory-Based Policy Learning
-Before presenting our main algorithm, Ada-NAV, we start by highlighting an interesting connection between a policy entropy and the spectral gap of its induced transition matrix.
+In the following section, we leverage this theorem to empirically establish a useful connection between policy entropy and spectral gap to control the exploration with no need to estimate $\mathbb{P}_{\pi_{\theta}}$.
+III Proposed Approach: Controlled Exploration-based Policy Learning
 III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
-Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap from the mixing time lower bound in Theorem 4.
+Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap $1-\lambda_{\theta}^{2}$ that lower bounds $\tau_{mix}^{\theta}$.
 However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$.
 One possible workaround is to estimate the transition kernel, but this quickly becomes intractable due to the $\Omega(|\mathcal{S}||\mathcal{A}|)$ number of samples required for each estimate.
-Interestingly, one key insight we observe is that if the policy becomes more deterministic and thus the entropy decreases, the spectral gap of the induced transition matrix also decreases or stays the same, suggesting a monotonic relationship between mixing time and policy determinism.
-Figure 2 illustrates this correlation with a 2D gridworld environment.
-We have a $25$-by-$25$ gridworld.
-Each plot in Figure 2 corresponds to a different fixed transition probability kernel $\mathbb{P}$.
-For each fixed kernel, we construct 100 tabular policies, each with a discrete action space of 5 elements to represent up, down, left, right, and stay.
-The tabular policies have various average entropies, ranging from maximum entropy with uniform probability over the action space for each state (end of x-axis on Figure 2) to completely deterministic for each state (beginning of x-axis) with a probability of $1$ of moving upwards.
-We define policy entropy in for tabular policies as the average Shannon entropy across all states, see Equation 5.
-For each policy, we multiply the policy with the fixed transition kernel to calculate the underlying Markov Chain and then estimate its spectral gap.
-In this environment, if the agent chooses an action that would make it go off the edge, the agent will stay in the current state.
-Both transition kernels are deterministic: each state-action pair is mapped to a single next state a potential agent can occupy.
-One kernel represents an empty gridworld with no obstacles.
-The relationship between spectral gap and policy entropy is plotted in orange for this kernel.
-The other kernel is represents a gridworld with $4$ walls dividing it into $4$ equal regions.
-Please see Appendix for visualization.
-The plot corresponding to this kernel is blue in Figure 2.
-[^1] Code to notebook to generate plot: https://github.com/Bridge00/spectral_gap_policy_entropy.git
-To be clear, in these experiments, we have not trained any agent to navigate in the gridworld.
-We have a static environment where we manually designed transition kernel matrices and multiplied each one with 100 different tabular policies to estimate the underlying Markov Chains and their corresponding spectral gaps, a task normally intractable for more complex, realistic environments.
-For results displaying Ada-NAV’s effectiveness in increasing sample efficiency for a goal-reaching task in this environment and other gridworld environments, please see the Appendix.
-Figure 2: Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over a $25$-by-$25$ gridworld environment with different fixed transition kernels.
+Interestingly, one key insight we observe in $25$-by-$25$ gridworld environments is that if the policy entropy decreases, the spectral gap of the induced transition matrix also decreases or stays the same (Figure 2).
+Remark: We have not trained any agent to navigate in the gridworld.
+We have a static environment where we manually designed transition kernel matrices and multiplied each one with $100$ different tabular policies with various entropies, from uniform distributions to deterministic, distributions to estimate the underlying Markov Chains and their corresponding spectral gaps.
+Because of the impracticality of calculating spectral gaps from more complex environments, we provide these proof-of-concept experiments with gridworld environments to support the formulation of our CCE approach.
+Figure 2: [LEFT] Visualization of $25$-by-$25$ gridworld with $4$ walls (black) dividing it into equal regions.
+[RIGHT] Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over two different gridworld environments.
 Policy entropy is a viable proxy for mixing time.
@@
 where $N$ is the action space dimension and $\Sigma_{\theta}(s)$ is the covariance matrix of the policy $\pi_{\theta}(\cdot|s)$.
-We note that there are no analytical equations to calculate mixing time, and calculating spectral gap requires knowledge of the transition dynamics $P(s^{\prime}|s,a)$ which restricts it to only model-based RL.
-Thus, using analytical Equations (5) and (6) as a proxy for mixing time is a more simple method that can be used for either model-free or model-based RL.
-III-B Ada-NAV: Mapping Policy Entropy to Trajectory Length
-We now present our novel adaptive trajectory length scheme for robotic navigation.
-The key to our approach is to select trajectory length based on the current policy entropy value.
-We perform this computation using a monotonic mapping from the range of possible policy entropy values to a range of possible trajectory length values.
+We note that there are no analytical equations to calculate mixing time, and calculating spectral gap requires knowledge of the transition dynamics $P(s^{\prime}|s,a)$ in the form of a matrix, which is infeasible for complex environments.
+Thus, using analytical Equations (4) and (5) as a proxy for mixing time is a simple method that can be used for model-free or model-based RL.
+III-B CCE: Confidence-Controlled Exploration
+We now present our novel CCE approach for robotic navigation.
+The key to our approach is to select trajectory length based on the current value of the policy entropy, our measure for confidence.
+We perform this computation using a monotonic mapping $f$ from the range of possible policy entropy values to a range of possible trajectory length values.
+In this paper, we utilize a linear function; however, we note that $f$ can be implemented with any monotonic mapping such as an exponential function.
 Let $H_{i}$ denote the maximum possible policy entropy value and $H_{c}$ the policy entropy of the current policy.
 In practice, we choose $H_{i}=H(\pi_{\theta_{0}})$ to be the entropy of the initial policy, since policy entropy tends to decrease during training.
-Let $t_{i}\in\mathbb{N}$ denote the user-specified initial trajectory lengths and let $\eta>0$ denote a user-specified sensitivity parameter that controls the rate of change of length (see the Appendix for specific values of $t_{i},\eta$ used).
-For our experiments, we determine the trajectory length corresponding to the current policy entropy using the mapping $f$ below:
-$t_{c}=f^{H_{i}}_{t_{i},\eta}(H_{c})=\text{round}\left(1-\frac{H_{c}}{H_{i}}(1-\frac{t_{i}}{\eta})\right),$ (7)
-where the rounding operation is to ensure that $t_{c}$ is always an integer.
-In the Appendix, we utilize other monotonic transformation to map entropy to trajectory length in the form of an exponential function.
+Let $K_{i}\in\mathbb{N}$ denote the user-specified initial trajectory lengths and let $\alpha>0$ denote a user-specified sensitivity parameter that controls the rate of change of length.
+For our experiments, we determine the trajectory length, $K_{c}$, corresponding to $H_{c}$ using the mapping $f$ below:
+$K_{c}=f^{H_{i}}_{K_{i},\alpha}(H_{c})=\text{round}\left(\alpha-\frac{H_{c}}{H_{i}}(\alpha-K_{i})\right),$ (6)
+where the rounding operation is to ensure that $K_{c}$ is always an integer.
 Furthermore, in our experiments, we initialize our policies with a Gaussian distribution.
-To ensure, that $H_{c}$ never increases too much so that $t_{c}$ becomes 00, one can initialize with a uniform distribution so that $H_{i}$ is the maximum possible entropy.
-Algorithm 1 summarizes our Ada-NAV methodology for adaptive trajectory length PG methods.
+To ensure, that $H_{c}$ never increases too much so that $K_{c}$ becomes 00, one can initialize with a uniform distribution so that $H_{i}$ is the maximum possible entropy.
+Algorithm 1 summarizes our CCE methodology for adaptive trajectory length PG methods.
 Note that any on-policy PG method can used to carry out the policy gradient update in line 3.
@@
 During each episode, a gradient update occurs after every trajectory.
-The amount of samples in each trajectory is the trajectory length.
-The episode ends when goal is reached or when a set number of interactions have occurred.
+The episode ends when the goal is reached or when a user-set number $T$ of interactions has occurred.
 For clarification, given an episode, the trajectory length is fixed.
-At the start of the new episode is when the trajectory length is adjusted based on policy entropy.
-1:  Initialize: policy parameter $\theta_{0}$, policy entropy $H_{i}$, trajectory length $t_{i}$, current trajectory length $t_{c}=t_{i}$, maximum trajectory length $t_{d}$, number of episodes $E$, maximum number $K$ of iterations per episode, sensitivity parameter $\eta$, $t=0$
+$K_{c}$ is only adjusted based on the policy entropy at the start of the new episode.
+1:  Initialize: policy parameter $\theta_{0}$, policy entropy $H_{i}$, trajectory length $K_{i}$, current trajectory length $K_{c}=t_{i}$, number of episodes $E$, maximum number $T$ of iterations per episode, sensitivity parameter $\alpha$, update index $u=0$
 2:  for $e=0,\ldots,E-1$ do
-3:     calculate $H_{c}$ using 5 for discrete actions, 6 for continuous actions
-4:     calculate $t_{c}$ using 7
-5:     $k=0$
-6:     while $k<K$ do
+3:     calculate $H_{c}$ using 4 for discrete actions, 5 for continuous actions
+4:     calculate $K_{c}$ using 6
+5:     $t=0$
+6:     while $t<T$ do
 […]
-8:        Perform gradient update using $\mathcal{T}$ to obtain $\theta_{t+1}$
-9:        $k=k+t_{c}$
-10:        $t=t+1$
+8:        Perform gradient update using $\mathcal{T}$ to obtain $\theta_{u+1}$
+9:        $t=t+K_{c}$
+10:        $u=u+1$
 11:     end while
 12:  end for
-13:  Return: $\pi_{\theta_{t}}$
-Algorithm 1 Policy Learning with Ada-NAV
+13:  Return: $\pi_{\theta_{u}}$
+Algorithm 1 Policy Learning with CCE
 IV Experiments and Results
-This section provides a detailed empirical study and comparison of the proposed adaptive trajectory methods in Algorithm 1.
+This section provides a detailed empirical study and comparison of the proposed CCE method in Algorithm 1.
 Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator.
-For REINFORCE, PPO, and SAC, using Ada-NAV as the trajectory length scheme leads to a navigation policy with a higher cumulative reward return than using constant trajectory lengths.
-The drawn trajectories are sample representations of the robot’s odometry data collected during each navigation task.
-The trajectories pictured are representative of all three algorithms, since we found that each trajectory-length scheme leads to similar navigation behavior regardless of which RL algorithm is used.
-See Figures 4 and 6 for training details.
-Figure 4: [TOP] Learning curves during training and [BOTTOM] sample complexity savings during navigation tasks for the even terrain shown on the left of Figure 3.
+For REINFORCE, PPO, and SAC, using CCE as the trajectory length scheme leads to a navigation policy with a shorter path length return than baselines with limited samples.
+The drawn trajectories are sample representations of the robot’s odometry data collected.
+The trajectories pictured are representative of all $3$ algorithms since we found that each trajectory-length scheme leads to similar navigation regardless of the RL algorithm.
+See Figure 4 for training details.
+Figure 4: [TOP] Learning curves during training and [BOTTOM] percent of sample budget of $\sim 4.5e4$ samples during navigation tasks for the even terrain shown on the left of Figure 3.
 Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths.
-For each algorithm, Ada-NAV converges to a higher cumulative reward while expending less of the total budget of $\sim 4.5e4$ samples.
+For each algorithm, CCE converges to a higher cumulative reward while expending less of the total budget.
 We ran $8$ to $10$ independent replications for each of the $12$ experiments.
-Although fixed trajectory length 200 performs comparably to Ada-NAV in terms of sample complexity, it achieves lower reward in all cases.
-Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained using Ada-NAV vs.
+Although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
+Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained with REINFORCE using CCE vs.
 fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$.
-We observe that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training.
-In contrast, Ada-NAV policies successfully complete the navigation tasks after training with the same sample budget.
-Further, the experiments with a real Clearpath Husky robot validate that our method can be transferred onto real robotic systems without significant performance degradation.
+We observed that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training.
+In contrast, CCE policies successfully complete the tasks after training with the same limited sample budget.
+The experiments with a real Clearpath Husky robot validate that CCE can be transferred onto real robotic systems without significant performance degradation.
 The drawn trajectories are sample representations of the robot’s odometry data collected during each task.
 RL Algorithms.
-We choose three RL algorithms to integrate Ada-NAV into for improved sample efficiency: REINFORCE, PPO, and SAC.
-As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that Ada-NAV can improve the sample efficiency of various classes of policy gradient algorithms.
+We choose three RL algorithms to integrate CCE into for improved sample efficiency: REINFORCE, PPO, and SAC.
+As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that CCE can improve the sample efficiency of various classes of policy gradient algorithms.
+For PPO and SAC, we adaptively change the size of the replay buffer.
+For information on hyperparameters, such as network architecture, $K_{i}$, and $\alpha$, please see Table II in the Appendix 111we re-provide the link to the supplemental version for convenience https://arxiv.org/abs/2306.06192 of the supplemental version.
 Baselines.
-Since our proposed approach relies on adaptive trajectory length, for each algorithm we test Ada-NAV with, we compare it against the same algorithm with fixed trajectory lengths.
-We also compare with fixed trajectory lengths with max-entropy regularizer as that also encourages exploration.
+Since our proposed approach relies on adaptive trajectory length, for each algorithm we test CCE with, we compare it against the same algorithm with fixed trajectory lengths.
+We also compare with fixed trajectory lengths with max-entropy regularizer as that encourages exploration.
 We also incorporate a classical navigation algorithm Ego-graph [51], to highlight the navigation performance of our method under challenging terrain conditions.
-Ego-graph is a baseline method in [48] that utilizes robot-centric elevation data to obtain the actions that minimize the elevation gradient cost.
-Figure 6: [TOP] Learning curves during training and [BOTTOM] sample complexity savings during navigation tasks for the uneven terrain shown on the right of Figure 3.
-Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths.
-For each algorithm, Ada-NAV converges to a higher cumulative reward while expending less of the total budget of $\sim 4.5e4$ samples.
-We ran $8$ to $10$ independent replications for each of the 12 experiments.
-Although fixed trajectory length 200 performs comparably to Ada-NAV in terms of sample complexity, it achieves lower reward in all cases.
+Ego-graph is a baseline method in [51] that utilizes elevation data to obtain the actions that minimize the elevation gradient cost.
+Figure 6: Similar to Figure 4 but for uneven terrain shown on the right of Figure 3.
+CCE still consistently reaches the highest cumulative reward while saving the largest percentage of training data.
+As in Figure 4, although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
 Environments: We utilize two environments in a unity-based outdoor robot simulator.
-The policies for the outdoor robot simulator are based on the REINFORCE algorithm [27].
 Finally, we deploy navigation policies trained for a fixed sample budget using the robot simulator on a real Clearpath Husky robot.
 We present the details of the training environments below:
-Even and Uneven Terrain Navigation in Robotic Simulations: We use a Unity-based outdoor simulator with a Clearpath Husky robot model to train policies for two navigation tasks: 1.
+Simulated Even and Uneven Terrain Navigation: We use a Unity-based outdoor simulator with a Clearpath Husky robot model to train policies for two navigation tasks: 1.
 Goal Reaching, 2.
@@
 We used identical policy distribution parameters and neural network architectures with all four algorithms during training and evaluation.
-The policies are implemented using Pytorch in to train in the simulator equipped with a Clearpath Husky robot model, and a ROS Melodic platform.
-The training and simulations are executed on a workstation with an Intel Xeon 3.6 GHz CPU and an Nvidia Titan GPU.
+The policies are trained using Pytorch in the simulator equipped with a Clearpath Husky robot model, and a ROS Melodic platform.
+The training simulations are executed on a workstation with an Intel Xeon 3.6 GHz CPU and an Nvidia Titan GPU.
 Our agent is a differential drive robot with a two-dimensional continuous action space $a=(v,\omega)$ (i.e.
 linear and angular velocities).
-The state space varies for the two navigation tasks and is obtained in real time from the simulated robot’s odometry and IMU sensors.
-The state inputs used in the two scenarios are, The simulator includes even and uneven outdoor terrains and a differential drive robot model with a two-dimensional continuous action space $a=(v,\omega)$ (i.e.
-linear and angular velocities).
-We perform two navigation tasks: 1.
-Even terrain, 2.
-Uneven terrain navigation using the robot’s local sensor data as state observations.
-The state space varies for the two navigation tasks and is obtained in real-time from the simulated robot’s odometry and IMU sensors.
+The simulator includes even and uneven outdoor terrains and the state space varies for the two navigation tasks and is obtained in real-time from the simulated robot’s odometry and IMU sensors.
 The state inputs used in the two scenarios are,
-  • $d_{goal}\in[0,1]$ Normalized distance (w.r.t the initial straight-line distance) between the robot and its goal;
-  • $\alpha_{goal}\in[0,\pi]$ Angle between the robot’s heading direction and the goal;
-  • $(\phi,\theta)\in[0,\pi]$ Roll and pitch angle of the robot respectively.
+  • $d_{goal}\in[0,1]$- Normalized distance (w.r.t the initial straight-line distance) between the robot and its goal;
+  • $\alpha_{goal}\in[0,\pi]$- Angle between the robot’s heading direction and the goal;
+  • $(\phi,\theta)\in[0,\pi]$- Roll and pitch of robot, respectively.
 Inspired by [26], we define three reward functions to formulate the sparse rewards to train our navigation policies: Goal distance reward $r_{dist}$, Goal heading reward $r_{head}$, and Elevation reward $r_{elev}$ to formulate the sparse rewards to train our navigation policies.
@@
 Similarly, $r_{elev}$ is defined as $r_{elev}=\mathbb{I}_{\{|\phi|\geq\pi/6\}}\cup\mathbb{I}_{\{|\theta|\geq\pi/6\}}$.
-We observe that our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature[31].
-(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position are given at each episode.
-The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{goal}+r_{head}$.
+Our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature [41].
+(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position given at each episode.
+The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{dist}+r_{head}$.
 (b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment.
-The state space is $s=[d_{goal},\alpha_{goal},\phi,\theta]$.
+To account for this complexity, we input the roll and pitch of the angle of the robot into the state space, $s=[d_{goal},\alpha_{goal},\phi,\theta]$.
 The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs.
-The overall reward formulation $r_{uneven}=r_{goal}+r_{head}+r_{elev}$.
-(c) Real-world Experiments: The real and simulated testing were conducted in different outdoor scenarios.
-Particularly, the real-world environments are reasonably different in terms of terrain structure (i.e.
-elevation gradient) and surface properties (i.e grass and tiny gravel regions with different levels of friction) which are not included or modeled in the simulator (e.g.
-elevation gain in the simulator is up to $\sim 4m$, however, we restricted it to $\sim 1.5m$ elevation gain in the real world for robot’s safety).
-Our Clearpath Husky robot is equipped with a VLP16 LiDAR and a laptop with an Intel i9 CPU and an Nvidia RTX 2080 GPU.
+The reward is $r_{uneven}=r_{dist}+r_{head}-r_{elev}$.
 […]
@@
 Reinforce: traj. length = 300 65 18.75 -
-Ada-NAV(Ours) 84 11.62 -
+CCE (Ours) 84 11.62 -
 Uneven Ego-graph [51] 54 20.36 2.189
@@
 Reinforce: traj. length = 300 31 32.44 2.236
-Ada-NAV(Ours) 72 16.28 1.985
-Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks at a fixed sample budget.
+CCE (Ours) 72 16.28 1.985
+Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks with limited samples.
 SR denotes success rate, PL denotes average path length, and EC denotes elevation cost.
@@
 IV-A Evaluation Metrics
-We use the following metrics to evaluate the quality of the navigation paths generated with a fixed sample budget:
   • Success rate: The percentage of successful goal-reaching attempts without robot flip-overs (especially in uneven terrain) out of the total number of trials.
-  • Average path length: The average length of the path taken to reach a goal located 10 meters away from the robot’s starting position.
+  • Average path length: The average length of path taken to reach goal 10 meters away from the starting position.
   • Elevation cost: The total elevation gradient experienced by the robot during a trial (i.e., if $z_{r}$ is the vector that includes the gradient of the vertical motions of the robot along a path, the elevation cost is given by $||\nabla z_{r}||$).
 IV-B Results and Discussions
-Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, Ada-NAV achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths.
-Significantly, despite providing comparable returns to Ada-NAV, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig.
+Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, CCE achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths.
+Significantly, despite providing comparable returns to CCE, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig.
 4 and 6).
-Please see the Appendix for similar results in 2D gridworld environments with actor-critic algorithm[49].
+For the agents using CCE, in the even terrain simulations, all agents started with an initial trajectory length of $180$, and in the uneven terrain, $200$.
+All those agents end with a final length of $300$.
 Navigation Performance: We compare our method’s navigation performance qualitatively in Fig.
 5 and quantitatively in Table I.
-To highlight the importance of sample efficiency during training, we use policies trained for a fixed sample budget to perform navigation tasks in even and uneven simulated terrains.
-We observe that, except Ada-NAV, the policies cannot consistently complete the navigation tasks, resulting in significantly lower success rates.
+To highlight the importance of sample efficiency during training, we use policies trained up to episode $150$ for a limited sample budget of $\sim 4.5e4$ to perform navigation tasks in even and uneven simulated terrains.
+The policies, except for ones from CCE, cannot consistently complete the navigation tasks, resulting in significantly lower success rates.
 Even successful trials of these policies lead to significantly longer paths with several loops instead of heading toward the goal.
-In contrast, Ada-NAV results in relatively straight and consistent paths toward the goals.
+In contrast, CCE results in relatively straight and consistent paths toward the goals.
 We further observe that uneven terrain navigation is a comparatively challenging task even for traditional methods such as Ego-graph[51].
 Hence, the success rates are relatively low due to the robot flip-overs that occurred during trials.
-However, we observe that Ada-NAV leads to higher success rates even compared to baseline methods such as ego-graph.
-We further observe that Ada-NAV results in comparatively lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions.
+However, CCE leads to higher success rates even compared to baseline methods such as ego-graph.
+We further observe that CCE results in lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions.
 Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot.
-We observe that our Ada-NAV even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies demonstrate inconsistent navigation performance similar to the simulations.
+With CCE, even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies show inconsistent performance similar to the simulations.
+Remark: Given more training samples, the baseline methods could perform better in success rate and path length during test time given that Figures 4 and 6 show the many times they can reach similar cumulative return as CCE.
+However, we tested policies from earlier in training to highlight that CCE does not need as many samples.
 V Conclusions, Limitations, and Future Works
-In this work, we propose an adaptive trajectory scheme, called Ada-NAV, which encourages exploration and induces training sample efficiency for goal-reaching navigation in the sparse reward setting with RL-based methods.
+We propose CCE, which controls exploration and induces training sample efficiency for goal-reaching navigation in the sparse reward setting with RL-based methods by adaptively changing the trajectory length during training.
 We do so by empirically establishing a positive correlation between a policy’s entropy and the spectral gap of its induced Markov Chain.
-We show, through experiments in navigation tasks, that Ada-NAV induces more sample-efficient training than both fixed and random trajectory length schemes without the need for tuning.
-In experiments on real robots, Ada-NAV-trained policies generally achieve better success rates and shorter path lengths than policies trained using constant and random trajectory lengths for the same sample budget.
+We show that CCE induces more sample-efficient training than baselines.
+On real robots, CCE-trained policies generally achieve better success rates and shorter path lengths than ones trained with constant trajectory lengths and entropy regularization for the same sample budget.
 An important limitation of the present work is the restricted scope of our experimental investigation of the connection between policy entropy and the induced spectral gap.
-We considered only simple gridworld environments for the purposes of this paper, yet this important connection merits further investigation and validation on more sophisticated challenges, such as the MuJoCo suite of RL problems [50].
-During real-world experiments, we observe that all the navigation policies often lead to high motor vibrations due to dynamically infeasible velocities provided as the action.
-Hence, further investigation is required to generate dynamically feasible and smooth actions for a better sim-to-real transfer.
+Although we cannot explore the connection in complex environments due to spectral gap calculation impracticality, it can still merit investigation for simple environments with continuous state and action spaces.
+During real-world experiments, all the navigation policies often lead to high motor vibrations from infeasible velocities provided as the action.
+Hence, further investigation is required to generate dynamically feasible and smooth actions for better sim-to-real transfer.

### v8

@@
 However, traditional RL methods often hinge on the availability of a dense reward structure [5], which may rarely be available in complex, unpredictable environments [6, 7, 8, 9].
-Thus, reward function sparsity [10] frequently arise when applying RL methods to robotics problems, such as robotic navigation, due to their simple to implement and does not require expert knowledge to design like dense reward structures.
-However, this simplicity comes with a sparse feedback signal to learn from, thus leading to a limited exploration of the environment, and ultimately, to a suboptimal policy [11].
+Thus, reward function sparsity [10] frequently arise when applying RL methods to robotics problems, such as robotic navigation, due to its simplicity to implementation and lack of need for expert knowledge to design required dense reward structures.
+However, this simplicity comes with a sparse feedback signal to learn from, thus leading to a limited exploration of the environment, and to a premature exploitation of a suboptimal policy [11].
 Existing literature in RL for robotics in sparse rewards has utilized methods such as entropy regularization [12], reward shaping [13, 14, 15, 16], curiosity-driven exploration [5, 17].
@@
 Another major challenge of existing works comes from the sample inefficiency of exploration in sparse rewards.
-This issue is significant in robotic systems, due to the simulation-intensive nature of RL methods and the computational expense of realistic robotics simulation platforms [18].
+This issue is significant in robotic systems, due to the simulation-intensive nature of RL methods and the high computational computational associated with realistic robotics simulation platforms [18].
 Thus, RL training typically requires finding an optimal policy in a constrained sample budget.
-In this work, we diverge from previous exploration paradigms to respect the practicality of limited training samples.
-We do so by controlling exploration, rather than encouraging it based on a novel connection between the uncertainty, or confidence, of a policy and policy gradient estimation.
+In this work, we propose an exploration paradigm to respect the practicality of limited training samples.
+We achieve this by balancing exploration, rather than encouraging it through a novel connection between the parameterized policy’s confidence and its gradient estimation.
 Figure 1: Robot navigation using policies trained with the REINFORCE policy gradient algorithm using four trajectory length schemes.
@@
 Main Contributions: We present Confidence-Controlled Exploration (CCE), a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods.
-We introduce using the policy confidence, measured by the average conditional policy entropy over all states to control exploration.
-We focus on balancing exploration with the policy entropy by adaptively changing a salient feature in most RL methods: trajectory length.
-In policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training.
-Our novel perspective is that exploration can be adaptively controlled by the number of samples needed to estimate the gradient.
-For navigation, an interaction is any action the mobile robot takes and the resulting state in the environment and the reward it receives.
+Our method leverages policy confidence, measured by the average conditional policy entropy across all states, to regulate exploration.
+Specifically, we focus on balancing exploration with policy entropy by adaptively adjusting a key feature in most RL methods: trajectory length, defined as the number of samples the robot collects in the environment before each gradient update during training.
+For navigation tasks, a data sample refers to any action the mobile robot takes at a given state, the resulting state transition in the environment, and the corresponding reward.
 To design CCE, our novel approach connects the need for the policy to explore to the mixing time of the policy.
@@
 When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [19].
-When mixing time is short, on the other hand, ensuring trajectory lengths are proportional frees up effort for performing more rapid policy updates and environmental exploration.
+Conversely, when mixing time is short, keeping trajectory lengths proportional to it frees up resources for faster policy updates and more efficient exploration of the environment.
 Unfortunately, mixing time is typically difficult to estimate in complex environments [20, 21], so proxy methods for determining trajectory length are needed.
@@
 Therefore, an alternative is estimating a lower bound.
-As described in [49], one such lower bound of mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{2}_{\theta}$:
+One such quantity that can lower bound of mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, $1-\lambda^{\theta}_{2}$, where $\lambda^{\theta}_{2}$ is the second largest eigenvalue of $\mathbb{P}_{\pi_{\theta}}$, as stated in the theorem below [49]:
+Theorem II.1
+Let $\tau_{mix}^{\theta}$ be the mixing time induced by policy $\pi_{\theta}$. Then,
+$\tau_{mix}^{\theta}(\epsilon)\geq\mathcal{O}\left(\frac{1}{1-|\lambda^{\theta}_{2}|}\right).$ (4)
 When an estimate of $\mathbb{P}_{\pi_{\theta}}$ is available, this provides a method for bounding its mixing time.
@@
 III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
-Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap $1-\lambda_{\theta}^{2}$ that lower bounds $\tau_{mix}^{\theta}$.
+Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap $1-\lambda^{\theta}_{2}$ that lower bounds $\tau_{mix}^{\theta}$.
 However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$.
@@
 Because of the impracticality of calculating spectral gaps from more complex environments, we provide these proof-of-concept experiments with gridworld environments to support the formulation of our CCE approach.
-Figure 2: [LEFT] Visualization of $25$-by-$25$ gridworld with $4$ walls (black) dividing it into equal regions.
+Figure 2: [LEFT] Visualization of $25\times 25$ gridworld with $4$ walls (black) dividing it into equal regions.
 [RIGHT] Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over two different gridworld environments.
@@
 2:  for $e=0,\ldots,E-1$ do
-3:     calculate $H_{c}$ using 4 for discrete actions, 5 for continuous actions
-4:     calculate $K_{c}$ using 6
+3:     calculate $H_{c}$ using 5 for discrete actions, 6 for continuous actions
+4:     calculate $K_{c}$ using 7
 5:     $t=0$
@@
 IV Experiments and Results
-This section provides a detailed empirical study and comparison of the proposed CCE method in Algorithm 1.
 Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator.
-For REINFORCE, PPO, and SAC, using CCE as the trajectory length scheme leads to a navigation policy with a shorter path length return than baselines with limited samples.
+For REINFORCE, PPO, and SAC, using CCE as the trajectory length scheme for $150$ episodes leads to a policy with a shorter path length return than baselines with limited samples.
 The drawn trajectories are sample representations of the robot’s odometry data collected.
 The trajectories pictured are representative of all $3$ algorithms since we found that each trajectory-length scheme leads to similar navigation regardless of the RL algorithm.
-See Figure 4 for training details.
-Figure 4: [TOP] Learning curves during training and [BOTTOM] percent of sample budget of $\sim 4.5e4$ samples during navigation tasks for the even terrain shown on the left of Figure 3.
+See Figures 4 and 6 for training details.
+Figure 4: [TOP] Learning curves during training and [BOTTOM] percent of sample budget ($\sim 4.5e4$) saved during navigation tasks for the even terrain (left of Figure 3).
 Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths.
@@
 We ran $8$ to $10$ independent replications for each of the $12$ experiments.
-Although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
-Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained with REINFORCE using CCE vs.
+Although a fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing episodes to end.
+Figure 5: Sample navigation trajectories generated in simulated (a, b) and real-world (c, d) environments by policies trained with REINFORCE using CCE vs.
 fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$.
@@
 For PPO and SAC, we adaptively change the size of the replay buffer.
-For information on hyperparameters, such as network architecture, $K_{i}$, and $\alpha$, please see Table II in the Appendix 111we re-provide the link to the supplemental version for convenience https://arxiv.org/abs/2306.06192 of the supplemental version.
+For information on hyperparameters, such as network architecture, $K_{i}$, and $\alpha$, please see the Appendix.
 Baselines.
@@
 Ego-graph is a baseline method in [51] that utilizes elevation data to obtain the actions that minimize the elevation gradient cost.
-Figure 6: Similar to Figure 4 but for uneven terrain shown on the right of Figure 3.
-CCE still consistently reaches the highest cumulative reward while saving the largest percentage of training data.
-As in Figure 4, although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
+Figure 6: [TOP] Learning curves during training and [BOTTOM] percent of sample budget ($\sim 4.5e4$) saved during navigation tasks for the uneven terrain (right of Figure 3).
+Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths.
+For each algorithm, CCE converges to a higher cumulative reward while expending less of the total budget.
+We ran $8$ to $10$ independent replications for each of the $12$ experiments.
+Although a fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing episodes to end.
 Environments: We utilize two environments in a unity-based outdoor robot simulator.
@@
   • $d_{goal}\in[0,1]$- Normalized distance (w.r.t the initial straight-line distance) between the robot and its goal;
-  • $\alpha_{goal}\in[0,\pi]$- Angle between the robot’s heading direction and the goal;
+  • $\zeta_{goal}\in[0,\pi]$- Angle between the robot’s heading direction and the goal;
   • $(\phi,\theta)\in[0,\pi]$- Roll and pitch of robot, respectively.
 Inspired by [26], we define three reward functions to formulate the sparse rewards to train our navigation policies: Goal distance reward $r_{dist}$, Goal heading reward $r_{head}$, and Elevation reward $r_{elev}$ to formulate the sparse rewards to train our navigation policies.
-Hence,
-$r_{dist}=\frac{\beta}{2}\mathcal{N}(\frac{d_{goal}}{2},\sigma^{2})+\beta\mathcal{N}(0,\sigma^{2}),$ (7)
-$r_{head}=\mathbb{I}_{\{|\alpha_{goal}|\leq\pi/3\}},$ (8)
-where, $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter.
+Hence, $r_{dist}=\frac{\beta}{2}\mathcal{N}(\frac{d_{goal}}{2},\sigma^{2})+\beta\mathcal{N}(0,\sigma^{2})$ and $r_{head}=\mathbb{I}_{\{|\zeta_{goal}|\leq\pi/3\}}$, where $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter.
 We set the variance $\sigma=0.2$ to ensure that the $r_{dist}$ only includes two narrow reward peaks near the goal and half-way from the goal.
@@
 (a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position given at each episode.
-The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{dist}+r_{head}$.
+The state space is $s=[d_{goal},\zeta_{goal}]$ and the overall reward formulation $r_{even}=r_{dist}+r_{head}$.
 (b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment.
-To account for this complexity, we input the roll and pitch of the angle of the robot into the state space, $s=[d_{goal},\alpha_{goal},\phi,\theta]$.
+To account for this complexity, we input the roll and pitch of the angle of the robot into the state space, $s=[d_{goal},\zeta_{goal},\phi,\theta]$.
 The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs.
@@
 Each method is tested for 100 trials in the outdoor simulator and the average values are reported in the table.
+CCE has a higher SR, lower PL, and lower EC than baseline approaches in the even terrain, and higher SR and lower EC in uneven.
 Test environments are different from training environments.
-IV-A Evaluation Metrics
+IV-A Evaluation Metrics for Navigation Paths Generated with Fixed Training Sample Budget
   • Success rate: The percentage of successful goal-reaching attempts without robot flip-overs (especially in uneven terrain) out of the total number of trials.
@@
 IV-B Results and Discussions
-Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, CCE achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths.
+Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, CCE achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed trajectory lengths and entropy regularization.
 Significantly, despite providing comparable returns to CCE, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig.
@@
 Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot.
-With CCE, even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies show inconsistent performance similar to the simulations.
+With CCE, even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in success rate, path length, and elevation cost while other policies show inconsistent performance similar to the simulations.
 Remark: Given more training samples, the baseline methods could perform better in success rate and path length during test time given that Figures 4 and 6 show the many times they can reach similar cumulative return as CCE.
@@
 On real robots, CCE-trained policies generally achieve better success rates and shorter path lengths than ones trained with constant trajectory lengths and entropy regularization for the same sample budget.
-An important limitation of the present work is the restricted scope of our experimental investigation of the connection between policy entropy and the induced spectral gap.
+A limitation is the scope of our experimental investigation of the connection between policy entropy and the induced spectral gap.
 Although we cannot explore the connection in complex environments due to spectral gap calculation impracticality, it can still merit investigation for simple environments with continuous state and action spaces.
-During real-world experiments, all the navigation policies often lead to high motor vibrations from infeasible velocities provided as the action.
+In real-world experiments, policies often lead to high motor vibrations from infeasible velocities provided as actions.
 Hence, further investigation is required to generate dynamically feasible and smooth actions for better sim-to-real transfer.